OpenStudy (kainui):

Infinitely nested 2x2 matrix...

OpenStudy (kainui):

\[\left[\begin{matrix}\left[\begin{matrix}\left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right] \\ \left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right]\end{matrix}\right] & \left[\begin{matrix}\left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right] \\ \left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right]\end{matrix}\right] \\ \left[\begin{matrix}\left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right] \\ \left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right]\end{matrix}\right] & \left[\begin{matrix}\left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right] \\ \left[\begin{matrix} & \\ & \end{matrix}\right] & \left[\begin{matrix} & \\ & \end{matrix}\right]\end{matrix}\right]\end{matrix}\right]\] So it's equal to its own determinant...What's the matrix? lol ok maybe this is too simple, but I wonder if someone has any fun ideas to go with this.

Parth (parthkohli):

I wonder what happens if we try to multiply this by any scalar and try to write the result at the deepest level.

Parth (parthkohli):

Of course there's no deepest level, so is this matrix multiplied by a scalar equal to the matrix itself? :P

Parth (parthkohli):

lol, that was a joke. Good question by the way.

OpenStudy (kainui):

Well since the matrix seems to be the zero matrix (I think) I don't know if multiplying it by a scalar will even matter!

Parth (parthkohli):

Oh yeah, right. But don't take my replies seriously. Hehe.

OpenStudy (kainui):

Another fun to think about, what about having a matrix as the base of a logarithm? Is there some "natural" base like e is to scalars? Like, maybe the rotation matrix? \[\log_{\left[\begin{matrix}? & ? \\ ? & ?\end{matrix}\right]}(A)=B\]lol

OpenStudy (kainui):

Don't take my questions seriously! =P

OpenStudy (kainui):

\[A=\left[\begin{matrix}A & B \\ B & A\end{matrix}\right]\]\[B=\left[\begin{matrix}B & A \\ A & B\end{matrix}\right]\] Just a random thought, is this possible to be satisfied? If so, what are the matrices?